A 50.0 mL solution of H2SO4 has a concentration of 2.00 mol/L. It is dilutedby adding 150mL of water. What is the molar concentration of the sulphuric acid after dilution?


1
Expert's answer
2022-06-08T04:47:03-0400

With the dilution, the mole number remains constant:

"n_1=n_2"

"c_1V_1 = c_2V_2"

"c_2 = \frac{c_1V_1}{V_2} = \frac{2.00 \times 50.0}{150.0} = 0.67 mol/L"